OpenStudy (anonymous):

Given triangle ABC and side a = 21; side b = 20. Find the measurement of angle B if angle C is a right angle.

OpenStudy (jdoe0001):

 http://fouss.pbworks.com/f/tigratios.JPG \(\bf tan(\theta) = \cfrac{opposite}{adjacent}\implies tan^{-1}(tan(\theta) = tan^{-1}\left(\cfrac{opposite}{adjacent}\right)\\ \quad \\ \theta = tan^{-1}\left(\cfrac{opposite}{adjacent}\right)\)

OpenStudy (anonymous):

so would it be 46.4

OpenStudy (jdoe0001):

well.... I get .. something. else... make sure your calculator is in Degrees mode

OpenStudy (anonymous):

ok 17.8 would be correct then

OpenStudy (jdoe0001):

yw